OpenStudy (anonymous):

graph y=sinx and y=1/2sin2x using a table of values ( this is what i have done so far, the y=1/2sin2x doesn't seem right)

OpenStudy (anonymous):

pick values of \(x\) like \(\frac{\pi}{8},\frac{\pi}{4},\frac{3\pi}{8},\frac{\pi}{2}\) etc

OpenStudy (anonymous):